An improvised billiard ball is used in a local play where a 35 gram billiard ball resting in a corner is suddenly hit by another 45 grams ball moving 0.5 m/s. After the collision, the 45 grams ball remains at rest while the other ball moves away. Find the velocity of the ball moving away ? 


1
Expert's answer
2021-05-12T12:20:18-0400

Assuming it was a central collision and the 35-gram ball moves in a straight line, we can apply linear momentum conservation:


"p_i=p_f,\\
mv_m+Mv_M=mu_m+Mu_M,\\
0+Mv_M=mu_m+0,\\
Mv_M=mu_m,\\
u_m=v_M\frac Mm=0.5\frac{45}{35}=0.64\text{ m/s}."